Direct Materials and Direct Labor Variances

At the beginning of June, Kimber Toy Company budgeted 19,000 toy action figures to be manufactured in June at standard direct materials and direct labor costs as follows:

Direct materials $33,250
Direct labor 13,300
Total $46,550

The standard materials price is $0.70 per pound. The standard direct labor rate is $14.00 per hour. At the end of June, the actual direct materials and direct labor costs were as follows:

Actual direct materials $31,200
Actual direct labor 12,500
Total $43,700

There were no direct materials price or direct labor rate variances for June. In addition, assume no changes in the direct materials inventory balances in June. Kimber Toy Company actually produced 17,300 units during June.

Determine the direct materials quantity and direct labor time variances. Round your per unit computations to two decimal places, if required. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Direct materials quantity variance $
Direct labor time variance $
Answer #1

When direct material price and direct labor rate variance does not exist so total variance is equal to direct material quantity and direct labor time variance

Direct material quantity variance = (33250/19000*17300)-31200 = 925 Unfavorable

Direct labor time variance = (13300/19000*17300)-12500 = 390 Unfavorable